1 hose can fill a goldfish pond in 65 mins, and 2 hoses can the same pond in 40 mins. find how long it takes the second hose alone to fill the goldfish pond
Answer by nerdybill(7384)

1 hose can fill a goldfish pond in 65 mins, and 2 hoses can the same pond in 40 mins. find how long it takes the second hose alone to fill the goldfish pond
.
Let x = time (minutes) it takes for second hose alone
then
40(1/65 + 1/x) = 1
multiply both sides by 65x:
40(x + 65) = 65x
40x + 2600 = 65x
2600 = 25x
104 minutes = x